OpenStudy (anonymous):

Find the x-intercept and y-intercept of: 6x-8y=2 I need a bit of explanation and help in understanding this!

OpenStudy (anonymous):

to find the y intercept you need to express in this form: y=mx+c

OpenStudy (anonymous):

Alright so how do you use that formula to plug in the values?

OpenStudy (anonymous):

m is the gradient and c is the y intercept can you express in this form for the equation goven?

OpenStudy (anonymous):

I'm not sure but I'll give it a shot

OpenStudy (anonymous):

try and I will guide you through it

OpenStudy (anonymous):

hahah nevermind I'm lost

OpenStudy (anonymous):

6x-8y=2 now subtract 6x from both sides to get -8y=-6x+2 ok so far?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Now divide both sides by -8 to get: \[y=\frac{ 3 }{ 4 }x-\frac{ 1 }{ 4 }\]ok?

OpenStudy (anonymous):

okay that's slightly confusing but i get it

OpenStudy (anonymous):

Compare to y=mx+c above, what is the y intercept?

OpenStudy (anonymous):

-1/4 ?

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

to find the x intercept, you put y=0 into either the original equation or the second equation and find what is the value of x. Let me know if you need help.

OpenStudy (anonymous):

wait so do you end up dividing 3/4 by 1/4 ?

OpenStudy (anonymous):

The other way round. You divide 1/4 by 3/4

OpenStudy (anonymous):

okay so my answer would be 1/3 ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so the x intercept would be 1/3 and the y intercept would be -1/4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

Welcomed